Question 7 What is the value (in V) of Eºcell for the following reaction? Zn2+ (aq) + Pb (s) → Zn (s) + Pb2+ (aq)
Answer #1

For the given cell reaction

Cathode reaction: Zn2+ (aq) +  2e \to Zn (s) ; EoZn2+/Zn = - 0.76 V.

Anode reaction: Pb (s) - 2e \to Pb2+ (aq)  ; EoPb2+/Pb = - 0.13 V

Eocell = Eocathode - Eoanode = - 0.76 - (-0.13) = - 0.63 V.